# ReaDy-Go: Real-to-Sim Dynamic 3D Gaussian Splatting Simulation for Environment-Specific Visual Navigation with Moving Obstacles

### v1

@@ IV-E Real-World Visual Navigation Performance
To investigate whether the proposed pipeline improves visual navigation policies by mitigating the sim-to-real gap, we compare the real-world navigation performance of ReaDy-Go and baselines in the target environments, as illustrated in Fig. 5. For baselines in real-world experiments, Vid2Sim and ViNT are used, as they achieve better performance among the four baseline models in the simulation experiments.
We observe three main messages from the results in Table II. First, the photorealistic real-to-sim ReaDy-Go navigation simulation facilitates sim-to-real transfer of visual navigation policies trained solely in simulation. ReaDy-Go achieves comparable success rates in both Static and Dynamic tasks in the real world consistent with its simulation results across all environments, even though the real-world test environments exhibit minor appearance changes between data collection and deployment (e.g., floor texture, lighting, or background variations). Photorealistic datasets for dynamic scenarios improve the robustness of end-to-end visual navigation policies to the sim-to-real gap.
Second, environment-specific policies (ReaDy-Go and Vid2Sim) trained on real-to-sim datasets for target environments achieve higher success rates and lower average reaching times than the general navigation model (ViNT), although ViNT requires $15\times$ more parameters. This makes ReaDy-Go a practical pipeline for robots operating in specific environments, such as households, restaurants, and factories.
Lastly, photorealistic dynamic obstacles in ReaDy-Go appear to be important for maintaining navigation performance in dynamic environments. In Dynamic, ReaDy-Go shows the highest success rate and the lowest average reaching time, with only a slight performance degradation compared to Static. In contrast, Vid2Sim exhibits a larger performance drop in Dynamic compared to ReaDy-Go.
@@ IV-F Zero-Shot Sim-to-Real Generalization
We further validate the generalization potential of ReaDy-Go by measuring zero-shot sim-to-real performance in an unseen environment. For this experiment, the policy is trained on the combined datasets from three environments, i.e., a total of 1,200 episodes from Outside, Lobby, and Library. We then deploy the policy to the unseen environment, as shown in Fig. 5–(d). The policy achieves over a 50% success rate in both Static and Dynamic tasks, with a higher average reaching time compared to the results in the training environments, as depicted in Table III. These results imply that the policy learns general navigation behaviors, such as detouring around static and dynamic obstacles while reaching the goal, but requires a longer time to complete the task, as the policy does not have prior training in the unseen scene. ReaDy-Go could serve as a scalable approach toward general navigation if trained on substantially more navigation data from diverse environments, which can be easily reconstructed from a few minutes of videos using the proposed photorealistic real-to-sim dynamic GS simulation pipeline.
@@ V Conclusion
In this work, we propose ReaDy-Go, a real-to-sim dynamic GS simulation pipeline for training visual navigation policies. It generates photorealistic navigation datasets for dynamic scenarios in target environments by combining a GS scene, a human animation module, an expert planner, and a human planner. The navigation policy trained on the datasets achieves better navigation performance in target environments compared to baseline methods, including a prior real-to-sim simulation method and general navigation models, in both simulation and the real world. In other words, ReaDy-Go facilitates robust sim-to-real transfer even in dynamic environments and highlights the advantages of environment-specific policies and photorealistic dynamic obstacles. Furthermore, ReaDy-Go shows generalization through zero-shot sim-to-real deployment in an unseen environment. For future work, ReaDy-Go can be advanced in its various components, such as increasing the number of training environments to improve generalization or integrating reinforcement learning to further enhance navigation policies.

### v2

@@ IV-E Real-World Visual Navigation Performance
To investigate the effectiveness of the proposed pipeline in mitigating the sim-to-real gap, we compare the real-world navigation performance of ReaDy-Go and baselines in the target environments shown in Fig. 5. For baselines in real-world experiments, Vid2Sim and ViNT are used, as they achieve better performance among the four baseline models in the simulation experiments.
Table II reports three main findings. First, the photorealistic real-to-sim simulation of ReaDy-Go facilitates sim-to-real transfer of visual navigation policies trained solely in simulation. ReaDy-Go achieves comparable success rates in both Static and Dynamic tasks in the real world, consistent with its simulation results across all environments, even though the real-world test environments exhibit minor appearance changes between data collection and deployment (e.g., floor texture, lighting, or background variations). This indicates that photorealistic datasets for dynamic scenarios improve the robustness of end-to-end visual navigation policies to the sim-to-real gap.
Second, environment-specific policies (ReaDy-Go and Vid2Sim) trained on real-to-sim datasets for target environments achieve higher success rates and lower average reaching times than the general navigation model (ViNT), although ViNT requires $15\times$ more parameters. This validates the practicality of ReaDy-Go for robots operating in specific environments, such as households, restaurants, and factories.
Lastly, photorealistic dynamic obstacles in ReaDy-Go are important for maintaining navigation performance in dynamic environments. In Dynamic, ReaDy-Go shows the highest success rate and the lowest average reaching time, with only a slight performance degradation compared to Static. In contrast, Vid2Sim exhibits a larger performance drop in Dynamic compared to ReaDy-Go.
@@ IV-F Zero-Shot Sim-to-Real Generalization
We further assess the generalization potential of ReaDy-Go by measuring zero-shot sim-to-real performance in an unseen environment. For this experiment, the policy is trained on the combined datasets from three environments, i.e., a total of 1,200 episodes from Outside, Lobby, and Library. We then deploy the policy to an unseen environment, as shown in Fig. 5–(d). The policy achieves over a 50% success rate in both Static and Dynamic tasks, with the higher average reaching time compared to the results in the training environments, as depicted in Table III. These results suggest that the policy learns general navigation behaviors, such as detouring around static and dynamic obstacles while reaching the goal, but requires a longer time to complete the task, as the policy does not have prior training in the unseen scene. ReaDy-Go can provide a scalable approach toward general navigation if trained on larger navigation datasets from diverse environments, which can be easily reconstructed from a few minutes of videos using the proposed photorealistic real-to-sim dynamic GS simulation pipeline.
@@ V Conclusion
In this work, we propose ReaDy-Go, a real-to-sim dynamic GS simulation pipeline for training visual navigation policies. By integrating a GS scene, a human animation module, an expert planner, and a human planner, it generates photorealistic navigation datasets for dynamic scenarios in target environments. The navigation policy trained on the datasets achieves better navigation performance in target environments compared to baseline methods, including a prior real-to-sim simulation method and general navigation models, in both simulation and the real world. The results confirm that ReaDy-Go facilitates robust sim-to-real transfer even in dynamic environments and highlights the advantages of environment-specific policies and photorealistic dynamic obstacles. Furthermore, ReaDy-Go shows generalization through zero-shot sim-to-real deployment in an unseen environment. Future work includes expanding training environments to strengthen generalization or incorporating reinforcement learning to enhance navigation policies.